# Graph-Reward-SQL: Execution-Free Reinforcement Learning for Text-to-SQL via Graph Matching and Stepwise Reward

### v2

@@ 9 Limitations
While our proposed reward models demonstrate strong performance in text-to-SQL tasks and show much better performance for execution-based reward model, its SQL-specific architectural design currently restricts its applicability to broader code generation domains. Generalizing our framework to support languages such as Python or C++ would require some modifications to the program language representation design mechanisms, particularly to address the challenge of assessing functional equivalence across diverse programming paradigms.
However, this limitation opens promising directions for future work. We aim to explore the applicability of the execution-free reward model in broader code generation tasks, thereby offering viable alternatives to the common RL practice of relying solely on execution outcomes. This extension may contribute to the development of more general, efficient and flexible reward models for more code generation tasks.
@@ Appendix A Datasets Details
(Li et al., 2024b) is another large-scale and cross-domain Text-to-SQL dataset known for its complexity. It contains 12,751 question-SQL pairs drawn from 95 databases across more than 37 professional fields, making it a challenging testbed for evaluating the generalization capability of semantic parsers. A notable characteristic of BIRD is its “dirty” nature: queries may be inaccurate, column names are often ambiguous or poorly described, and databases may contain null values or irregular encodings. These issues collectively pose unique challenges for model robustness. The dataset is split into 9,428 training examples and 1,534 development examples, with the remaining held out for testing.
@@
comprises 200,000 examples related to the text-to-SQL task. Each data entry also includes a question, a database schema, and a reference SQL statement. To prepare the model for generating clean SQL queries during the PPO phase, we first fine-tune the Deepseek-coder-1.3b-ins model via Supervised Fine-Tuning (SFT), which enables the model to follow instructions and complete Text-to-SQL tasks effectively. The prompt used is detailed in O. Dataset is available at: https://huggingface.co/datasets/philikai/200k-Text2SQL.
@@ Appendix B Evaluation Setting Details
TS Performance provides a more robust and reliable assessment of semantic accuracy. For example, the original academic.sqlite database is automatically augmented into 411 variants, resulting in a total of 412 test databases. Each predicted-reference SQL pair is executed on all of them. This significantly reduces the likelihood of false positives and strengthens the evaluation’s ability to detect semantic discrepancies.
In our evaluation using the official TS framework, we emphasize strictness in semantic verification. Specifically, we enable the --keep_distinct flag to preserve the use of the DISTINCT keyword during evaluation. Meanwhile, we disable --plug_value to avoid artificially injecting ground-truth values into predicted queries, thus requiring the model to generate complete SQL outputs, including correct values. We use the TS code to evaluate Spider on its augmented dev split databases. For BIRD, since no augmented databases are provided, we apply the TS code directly on its original database. Code and Augmented databases are available at: https://github.com/taoyds/test-suite-sql-eval.
@@ Appendix C Implement Details
All experiments are conducted on an edge server equipped with NVIDIA A100-SXM GPUs. We extend the DeepSpeed-Chat framework (Yao et al., 2023) to adapt different reward models for Text-to-SQL tasks. During SFT, we use the AdamW optimizer with an initial learning rate of 5e-6. The batch size is 4. A polynomial learning rate scheduler is adopted to gradually decay the learning rate throughout training.
@@ C.1 Outcome Reward Experiments
For the outcome reward PPO experiments, we first conduct a warm-up phase using a different dataset from the target training corpus—200k-Text2SQL. We sample an equivalent number of examples as in Spider-train and perform 2 epochs of SFT warm-up using SFT. The sampling process also involves removing irrelevant samples, such as those with queries in Chinese, to ensure consistency and relevance in the training data. This warmup helps the model adapt to prompt-following behavior and prevents it from generating non-SQL artifacts. After SFT warmup, PPO training proceeds on the Spider/BIRD dataset with hyperparameters shown in table 5.
@@ C.2 Stepwise Reward Experiments
For the stepwise reward PPO experiments, we encourage the generation of CTE-structured SQLs by incorporating CTE-rewritten data into the SFT stage. Since SQL queries in BIRD are generally more complex than those in Spider, they are particularly well-suited for decomposition into modular subqueries using the CTE format. To leverage this, we rewrite the reference SQLs in the BIRD training set into CTE format using GPT, following the prompt strategies outlined in Appendix O. The resulting data is then combined with the warm-up dataset used in the outcome reward model experiments to construct a SFT warm-up set. This setup encourages the model to generate CTE SQL, enabling effective application of stepwise reward signals during PPO training.
@@
We applied reinforcement learning directly to the Qwen2.5-Coder-7B-Ins model using two distinct RL protocols: GRPO and PPO. Unlike PPO experiments conducted with Deepseek-Coder-1.3b-ins and Deepseek-Coder-6.7b-ins, we bypassed SFT and instead adopted a cold-start approach.
In experiments of Group Relative Policy Optimization (GRPO) with Qwen2.5-Coder-7B-Ins, we configured the learning rates for the actor and critic to 5e-6 and 2e-6, respectively, with a warm-up period of 10 steps to ensure stable training dynamics in the early stages. The model was trained with a batch size of 512, and mini-batches of size 256 were used. A weight decay of 0.1 was applied to regularize the model and mitigate overfitting. The group size for GRPO was set to 16, which corresponds to the number of completions generated per input prompt, thereby facilitating the optimization of the policy through relative comparisons between groups. For effective handling of long sequences, we set the maximum prompt length to 4096 and the maximum response length to 16,384. The training process was distributed across 8 machines, each equipped with 4 NVIDIA A100 80GB GPUs
@@
Figure 5: Reward score trends on the training set during GRPO training. The green curve corresponds to the execution-based reward (EX), while the orange curve shows the results using the proposed GMNScore. Both models are trained with the Qwen2.5-Coder-7B-Ins backbone under the GRPO framework. While GMNScore starts from a lower initial score compared to EX, it demonstrates a larger overall increase.
@@ Appendix D GRPO
In addition to experiments based on PPO, we also evaluate our methods under the Group Relative Policy Optimization (GRPO) framework Shao et al. (2024), in order to further assess the impact of different reward designs. In the standard text-to-SQL setting, $x$ denotes a natural language query, and $\mathcal{G}=\{\hat{q}_{1},\dots,\hat{q}_{G}\}$ denotes a group of $G$ candidate SQL queries sampled from the current policy $\pi_{\theta}(\cdot|x)$. To better leverage relative preferences among candidates, GRPO optimizes the policy by maximizing the following objective:
@@
[…]
@@
Appendix E Details of Execution Accuracy (EX)
Execution Accuracy (EX) is a commonly used reward signal in text-to-SQL tasks, offering supervision based on the correctness of query execution. However, we do not adopt the naive binary EX metric that only distinguishes between success and failure. Instead, we adopt a stronger variant of execution-based reward that integrates syntax check signals and runtime diagnostics, resulting in a more fine-grained supervision signal. This reward formulation—motivated by Pourreza et al. (2025b); Shojaee et al. (2023)—can be viewed as a syntax-aware extension of traditional EX, and is used as our primary execution-based baseline for comparison.
@@ Appendix E Details of Execution Accuracy (EX)
For the false positive problem of execution-based rewards, Test Suite (TS) (Zhong et al., 2020) attempted to improve robustness by using a set of test cases to simulate query behavior under different data distributions. Nevertheless, as shown in Zhan et al. (2025), both false positives (semantically incorrect queries that return the right result) and false negatives (semantically correct queries that fail under test data) persist, due to the reliance on incomplete or ambiguous database contents.
@@
Appendix F Cost Analysis of Reward Models
To evaluate the computational efficiency of different reward models, we conduct experiments on an edge server equipped with an Intel Xeon Platinum 8336C CPU (128 cores), 8 × NVIDIA A100-SXM GPUs, and a total memory capacity of 2.0 TiB. Our comparison focuses exclusively on model-based reward functions. Rule-based reward methods (AstPM, RelPM and our proposed StepRTM) are excluded from this analysis as they incur negligible memory overhead and execute efficiently without requiring GPU resources.
@@ Appendix F Cost Analysis of Reward Models
We use the training split of the Spider dataset and perform one epoch of PPO training, during which we measure the total time consumed by reward score computation. We then record the average reward execution time per sample, offering insights into the computational efficiency of the different reward models. Both BTRM and GMNScore are implemented using bfloat16 precision for acceleration. Execution-based reward (EX) incurs the highest computational cost due to repeated database calls. Note that, while EX can leverage high concurrency when executed on the CPU to accelerate performance, it still remains slower in practice due to its reliance on database calls. In real-world scenarios, different GPU rollouts of samples are processed concurrently through the reward model, which can improve processing speed. However, both BTRM and GMNScore can also easily achieve parallelization by loading the same model onto each GPU. Therefore, the speed reflected in the table is representative of the actual performance in training.
@@
sum(p.numel() for p in model.parameters())
We train Bradley-Terry Reward Model (BTRM) using DeepSeek-Coder-1.3B-Ins as base model, with a total of 1,346,471,936 parameters. In contrast, GMNScore adopts a lightweight Graph Matching Network with only 3,994,944 parameters.
@@
Appendix G Hybrid Outcome Reward Designs
@@
+ PPO w/ GMN & RelPM 66.34 16.69 41.52
+ PPO w/ EX & GMN 67.02 17.73 42.38
DeepSeek-Coder-6.7B-Ins 44.97 18.38 31.68
+ PPO w/ EX & RelPM 72.53 25.55 49.04
+ PPO w/ GMN & RelPM 72.34 26.40 49.37
+ PPO w/ EX & GMN 72.73 25.95 49.34
Table 7: TS Performance of DeepSeek-Coder-1.3B/6.7B-Ins models under different combinations of outcome reward strategies with Spider-Train.
Combining multiple outcome reward signals can further improve performance when their strengths complement each other. Table 8 shows that fusing EX with RelPM yields the highest average on Spider-Train (43.53% vs. 41.90% for GMN Score alone), while combining GMN Score with RelPM attains the top average on BIRD-Train (36.28%). By contrast, the EX+GMN Score combination yields less improvements or even declines, achieving only 42.38% on Spider and 33.33% on BIRD. We attribute this to both reward models emphasizing global correctness, resulting in feedback that is redundant and sometimes contradictory. The integration of RelPM’s localized partial matching feedback produces the most substantial improvements, highlighting the value of heterogeneous reward fusion that balances global execution accuracy with fine-grained structural supervision.
It is worth noting, however, that the effectiveness of combining multiple outcome rewards is significantly lower than that of our proposed Stepwise Reward model, StepRTM. The highest value achieved by the outcome rewards combination is 42.38%, whereas StepRTM achieves a higher value of 45.32%, representing an improvement of 2.94%. This highlights the effectiveness of our designed stepwise reward model StepRTM.
Appendix H Generalization of GMNScore Reward
@@
Although GMNScore Reward is initially trained solely on the Spider dataset, it generalizes across databases without requiring any additional fine-tuning. As shown in Table 8, when the policy is fine-tuned exclusively on BIRD-Train data, GMNScore still outperforms the EX reward in average Test-Suite Accuracy, consistent with its performance on Spider-Train.
@@
Appendix I Detailed Effectiveness of GMNScore Across GRPO
@@
+ GRPO w/ GMNScore (Ours) 78.53$\uparrow$2.13 29.92$\uparrow$1.43 54.23$\uparrow$1.78
@@
+ GRPO w/ GMNScore (Ours) 78.34$\uparrow$0.78 34.35$\uparrow$0.45 56.35$\uparrow$0.62
Table 9: TS Performance of Qwen2.5-Coder-7B/14B-Ins models directly trained by PPO/GRPO under EX and GMNScore outcome rewards with Spider-Train.
Our proposed GMNScore is not only effective with PPO, but also demonstrates effectiveness when applied to GRPO. Detail results are shown in the Table 9.
Appendix J StepRTM Calculation
@@
Appendix K Training Details of Neural Reward Model with Language odeling
@@
To train the reward model on above preference pairs, we finetune pretrained model DeepSeek-Coder-1.3B-Instruct with a reward head. Training is conducted using an effective batch size of $16\times 2$ (per-device batch size $\times$ gradient accumulation) with the AdamW optimizer, a learning rate of $1.41\times 10^{-5}$, and a linear warmup schedule. The model is trained for five epochs with a sequence length of 2048 tokens.
@@
The reward model should effectively evaluates generated SQL based on both syntactic clarity and execution-level correctness. To reinforcing schema understanding and functional equivalence, the data was structured as follows:
@@
Appendix L RelNode-based Partial Matching
@@
Based on RelNode representations, a partial matching strategy is employed to measure the similarity between the relational operator trees of the reference SQL $R(q^{\star})$ and the generated SQL $R(\hat{q})$. Let $\mathcal{N}_{q^{\star}}$ and $\mathcal{N}_{\hat{q}}$ denote the sets of nodes in the RelNodes of the reference and generated queries, respectively. A node pair $(n,n^{\prime})$ is considered a match, denoted as $\text{match}(n,n^{\prime})$, if the two nodes share the same relational operator type and their associated attributes are semantically equivalent.
@@
Precision =|$=\frac{|\{n\in\mathcal{N}_{\hat{q}}\mid\exists n^{\prime}\in\mathcal{N}_{q^{\star}},\ \text{match}(n,n^{\prime})\}|}{|\mathcal{N}_{\hat{q}}|}$
Recall =|$=\frac{|\{n\in\mathcal{N}_{q^{\star}}\mid\exists n^{\prime}\in\mathcal{N}_{\hat{q}},\ \text{match}(n,n^{\prime})\}|}{|\mathcal{N}_{q^{\star}}|}$
@@
Appendix N Prompt Strategy of CTE

### v3

@@ 9 Limitations
While our proposed reward models demonstrate strong performance in text-to-SQL tasks and show much better performance than the execution-based reward model, its SQL-specific architectural design currently restricts its applicability to broader code generation domains. Generalizing our reward model framework to support programming languages such as Python and Java would require some modifications to the language representation mechanisms, particularly to address the challenge of assessing functional equivalence across diverse programming paradigms.
However, this limitation opens promising directions for future work. We aim to explore the applicability of the execution-free reward model in broader code generation tasks in future work, thereby offering viable alternatives to the common RL practice of relying solely on execution outcomes. This extension may contribute to more general, efficient, and flexible reinforcement learning training for code generation tasks.
@@ Appendix A Datasets Details
(Li et al., 2024b) is another large-scale and cross-domain Text-to-SQL dataset known for its complexity. It contains 12,751 question-SQL pairs drawn from 95 databases across more than 37 professional fields, making it a challenging testbed for evaluating the generalization capability of semantic parsers. A notable characteristic of BIRD is its “dirty” nature: queries may be inaccurate, column names are often ambiguous or poorly described, and databases may contain null values or irregular encodings. These issues collectively pose unique challenges for model robustness. The dataset is split into 9,428 training examples and 1,534 development examples, with the remaining reserved for testing.
@@
comprises 200,000 examples related to the text-to-SQL task. Each data entry also includes a question, a database schema, and a reference SQL statement. To prepare the model for generating clean SQL queries during the PPO phase, we first fine-tune the Deepseek-coder-1.3b-ins model via Supervised Fine-Tuning (SFT), which enables the model to follow instructions and complete Text-to-SQL tasks effectively. The prompt used is detailed in Q. The dataset is available at: https://huggingface.co/datasets/philikai/200k-Text2SQL.
@@ Appendix B Evaluation Setting Details
Test-Suite (TS) Performance provides a robust and reliable assessment of semantic accuracy. For example, the original academic.sqlite database is automatically augmented into 411 variants, resulting in a total of 412 test databases. Each predicted-reference SQL pair is executed on all of them. This reduces the likelihood of false positives and strengthens the evaluation’s ability to detect semantic discrepancies.
In our evaluation using the official TS framework, we emphasize strictness in semantic verification. Specifically, we enable the --keep_distinct flag to preserve the use of the DISTINCT keyword during evaluation. Meanwhile, we disable --plug_value to avoid injecting ground-truth values into predicted queries, thus requiring the model to generate complete SQL outputs, including correct values. We use the TS code to evaluate Spider on its augmented dev split databases. For BIRD, since no augmented databases are provided, we apply the TS code directly on its original database. The code and augmented databases are available at: https://github.com/taoyds/test-suite-sql-eval.
In our experiments using the TS evaluation metric, all hints (e.g., age = year - birth_year) in BIRD dataset were deliberately omitted to align with the setup of the Spider dataset. This methodological choice results in lower performance metrics compared to studies that incorporated the hints available in the BIRD dataset.
@@ C.1 Outcome Reward Experiments
For the outcome reward PPO experiments, we first conduct a warm-up phase using a different dataset from the target training corpus—200k-Text2SQL. We sample an equivalent number of examples as in the training split of Spider dataset and perform 2 epochs of SFT warm-up using SFT. The sampling process also involves removing irrelevant samples, such as those with queries in Chinese, to ensure consistency and relevance in the training data. This warm-up helps the model adapt to prompt-following behavior and prevents it from generating non-SQL text. After SFT warm-up, PPO training proceeds on the Spider/BIRD dataset with hyperparameters shown in table 8. These hyperparameters, along with those for our reward model, were either adopted from prior work (Shojaee et al., 2023) or determined through empirical experiments.
@@ C.2 Stepwise Reward Experiments
For the stepwise reward PPO experiments, we encourage the generation of CTE-structured SQLs by incorporating CTE-rewritten data into the SFT stage. Since SQL queries in BIRD are generally more complex than those in Spider, they are particularly well-suited for decomposition into modular subqueries using the CTE format. To leverage this, we rewrite the reference SQLs in the BIRD training set into CTE format using GPT-4o, following the prompt strategies outlined in Appendix Q. The resulting data is then combined with the warm-up dataset used in the outcome reward model experiments to construct a SFT warm-up set. This setup encourages the model to generate CTE SQL, enabling effective application of stepwise reward signals during PPO training.
@@
Given the stronger capabilities of the base model, we directly applied reinforcement learning to the Qwen2.5-Coder-7B-Ins model using two distinct RL protocols, GRPO and PPO, without any warm-up training.
In experiments of Group Relative Policy Optimization (GRPO) with Qwen2.5-Coder-7B-Ins, we configured the learning rates for the actor and critic to 5e-6 and 2e-6, respectively, with a warm-up period of 10 steps to ensure stable training dynamics in the early stages. The model was trained with a batch size of 512, and mini-batches of size 256 were used. A weight decay of 0.1 was applied to regularize the model and mitigate overfitting. The group size for GRPO was set to 16, which corresponds to the number of completions generated per input prompt, thereby facilitating the optimization of the policy through relative comparisons between groups. For effective handling of long sequences, we set the maximum prompt length to 4096 and the maximum response length to 16,384.
@@
Figure 4: Reward score trends on the training set during GRPO training. The green curve corresponds to the execution-based reward model (EX), while the orange curve shows the results using the proposed reward model GMNScore. Both models are trained with the Qwen2.5-Coder-7B-Ins backbone under the GRPO framework. While GMNScore starts from a lower initial score compared to EX, it demonstrates a larger overall increase.
@@ Appendix D GRPO
In addition to experiments based on PPO, we also evaluate our methods under the Group Relative Policy Optimization (GRPO) RL framework Shao et al. (2024), in order to further assess the effectiveness of our proposed GMNScore outcome reward model. In the standard text-to-SQL setting, let $x$ denotes a natural language query, and $\mathcal{G}=\{\hat{q}_{1},\dots,\hat{q}_{G}\}$ denotes a group of $G$ candidate SQL queries sampled from the current policy $\pi_{\theta}(\cdot|x)$. To better leverage relative preferences among candidates, GRPO optimizes the policy by maximizing the following objective:
@@
$\bigg[$ $\frac{1}{G}\sum_{i=1}^{G}\min\left(\frac{\pi_{\theta}(\hat{q}_{i}|x)}{\pi_{\theta_{\mathrm{old}}}(\hat{q}_{i}|x)}A_{i},\;\right$.
@@
Appendix E Training of GMN of GMNscore
The Graph Matching Network (GMN) proposed serves as a foundation for assessing SQL query equivalence (Zhan et al., 2025), it was not originally designed for RL-based training. In our preliminary evaluation, reward model GMNScore built on the provided GMN model of FuncEvalGMN (Zhan et al., 2025) lagged behind that of the baseline EX under PPO experiments. We conducted a failure case analysis focusing on example where scores of GMNScore reward model disagreed with execution results. After manually checking, we identified several types of potential error situations.
To develop a GMN better suited for RL-based Text-to-SQL tasks, we retrain the model using an augmented version of the Spider-train-pair dataset introduced in (Zhan et al., 2025). The original Spider-train-pair dataset consists of 17,664 SQL query pairs annotated with binary equivalence labels. Two corresponding sets for evaluation, Spider-dev-pair and BIRD-dev-pair, contain 1,644 and 2,977 examples, respectively.
Seven augmentation strategies were crafted from the identified failure cases, adding 3,397 training pairs. The strategies are summarized in Table 9. After augmentation, the training set exceeded 20,000 SQL pairs.
Number Augmentation Type Original SQL Augmented Result
Enhancement of equivalent cases
137 IN Clause Replacement Original: SELECT Name FROM (SELECT Name, Age FROM technician) AS t
WHERE Age IN (36, 37)
Aug: SELECT Name FROM (SELECT Name, Age FROM technician) AS t
WHERE Age = 36 OR Age = 37
Enhancement of non-equivalent cases
422 Column Name Perturbation (Select Clause) Original: SELECT actid FROM activity
Aug: SELECT acti FROM activity
Aug: SELECT ctid FROM activity
566 Keyword Replacement (AND/OR) Original: SELECT * FROM Products
WHERE Price >= 60 AND Price <= 120
Aug: SELECT * FROM Products
WHERE Price >= 60 OR Price <= 120
1094 Symbol Replacement (Comparison Operator) Original: SELECT * FROM (SELECT dept_name, building FROM department) AS t
WHERE building > (SELECT AVG(building) FROM department)
Aug: SELECT * FROM (SELECT dept_name, building FROM department) AS t
WHERE building >= (SELECT AVG(building) FROM department)
276 Table Source Replacement Original: SELECT * FROM (SELECT name, email FROM user_profiles) AS t
WHERE name LIKE ’%Swift%’
Aug: SELECT * FROM (SELECT name, email FROM user) AS t
WHERE name LIKE ’%Swift%’
64 Column Name Replacement Original: SELECT candidate_id FROM candidate
ORDER BY candidate_id DESC LIMIT 3
Aug: SELECT people_id FROM candidate
ORDER BY people_id DESC LIMIT 3
838 Column Removal Original: SELECT circuitid, location FROM (SELECT circuitid, location, country FROM circuits) AS t
WHERE country = ’fraNce’ OR country = ’belGium’
Aug: SELECT circuitid FROM (SELECT circuitid, country FROM circuits) AS t
WHERE country = ’fraNce’ OR country = ’belGium’
Table 9: Examples of SQL augmentation strategies used for generating non-equivalent/equivalent SQL pairs. Each example introduces specific perturbations (e.g., column name change or logical operator replacement) that alter the semantics of the original SQL.
Appendix F Details of Execution Accuracy (EX)
Execution Accuracy (EX) is a commonly used reward signal in text-to-SQL tasks, offering supervision based on the correctness of query execution. However, we do not adopt the naive binary EX metric that only distinguishes between success and failure. Instead, we adopt a stronger variant of execution-based reward that integrates syntax check signals and runtime diagnostics, resulting in a more fine-grained supervision signal. Motivated by prior work Pourreza et al. (2025b); Shojaee et al. (2023), this reward formulation extends traditional binary execution accuracy with syntax awareness and serves as a strong execution-based baseline for comparison.
@@ Appendix F Details of Execution Accuracy (EX)
For the false positive problem of execution-based rewards, Test Suite (TS) (Zhong et al., 2020) attempted to improve robustness by using a set of test cases to simulate query behavior under different data distributions. Nevertheless, as shown in (Zhan et al., 2025), both false positives (semantically incorrect queries that return the right result) and false negatives (semantically correct queries that fail under test data) persist, due to the reliance on incomplete or ambiguous database contents.
@@
Appendix G Cost Analysis of Reward Models
To evaluate the computational efficiency of different reward models, we conduct experiments on an edge server equipped with an Intel Xeon Platinum 8336C CPU (128 cores) and a total memory capacity of 2.0 TiB. Our comparison focuses exclusively on model-based reward functions. Rule-based reward methods (AstPM/RelPM outcome reward model and our proposed stepwise reward model StepRTM) are excluded from this analysis as they incur negligible memory overhead.
@@ Appendix G Cost Analysis of Reward Models
We use the training split of the Spider dataset and perform one epoch of PPO training, during which we measure the total time consumed by the reward score computation. We then record the average reward calculation time per sample, providing insights into the computational efficiency of different reward models. Both BTRM and GMNScore are implemented using bfloat16 precision for acceleration. Execution-based reward (EX) incurs the highest computational cost due to repeated database calls. While EX can leverage high concurrency when executed on the CPU to accelerate performance, it still remains slower in practice due to its reliance on database calls. In real-world scenarios, different GPU rollouts of samples are processed concurrently through the reward model, which can improve processing speed. However, both BTRM and GMNScore can also easily achieve parallelization by loading the same model onto each GPU. Therefore, the speed reflected in the table is representative of the actual performance in training.
@@
p.numel() for p in model.parameters()
We train Bradley-Terry Reward Model (BTRM) using DeepSeek-Coder-1.3B-Ins as the base model, with a total of 1,346,471,936 parameters. In contrast, GMNScore adopts a lightweight Graph Matching Network with only 3,994,944 parameters.
@@
Appendix H Analysis of GMNScore Accuracy
Figure 6: There is a high Area Under the Curve (AUC) score between the GMNScore and execution results, consistently exceeding 97.6% during training.
Experimental results demonstrate the effectiveness of GMNScore as a reward model in PPO, significantly outperforming BTRM. We analyze the correlation [^5] AUC is used as a metric to assess the accuracy of similarity evaluation Zhan et al. (2025). In our experiments, a score of 100% does not indicate optimal performance, as false negatives are present in the execution outcomes. between these two reward signals and actual execution outcomes during PPO training. As shown in Figure 6, GMNScore consistently maintains a high correlation with the execution results. This indicates that GMNScore provides a more stable and precise reward signal than BTRM during training, contributing to its superior performance.
Appendix I Hybrid Outcome Reward Designs
@@
+ PPO w/ RelPM 62.86 14.67 38.77
+ PPO w/ EX 65.28 17.21 41.25
+ PPO w/ GMNScore 67.70 16.10 41.90
+ PPO w/ EX & AstPM 62.38 16.43 39.41
+ PPO w/ GMNScore & AstPM 65.28 16.17 40.73
+ PPO w/ GMNScore & RelPM 66.34 16.69 41.52
+ PPO w/ EX & GMNScore 67.02 17.73 42.38
@@
Table 11: TS Performance of DeepSeek-Coder-1.3B-Ins models under different combinations of outcome reward strategies.
We investigate whether combining multiple outcome reward models can further improve performance. Inspired by recent work that integrates multiple reward signals (Pourreza et al., 2025b; Ma et al., 2025), we explore various combinations among EX, GMNScore, RelPM, and AstPM. As shown in Table 11, combining EX with RelPM achieves the best overall performance, increasing the average accuracy from 41.90% (GMNScore alone) to 43.53%. The second-best result comes from EX combined with GMNScore, reaching 42.38%.
We attribute the relatively lower performance of EX & GMNScore compared to EX & RelPM to the fact that both EX and GMNScore emphasize global correctness. This may lead to redundant or even conflicting feedback, limiting the effectiveness of their combination. In contrast, the integration of RelPM provides localized partial matching signals, which offer complementary supervision and improve the performance.
Moreover, combining AstPM with either GMNScore or EX does not outperform using GMNScore or EX alone. We suspect this is due to AstPM’s focus on surface-level syntax. It may penalize syntactically different yet semantically equivalent SQLs, introducing false negatives that degrade training quality. As a result, this syntactic noise may undermine the semantic robustness provided by GMNScore or EX.
While GMNScore & RelPM performs better than RelPM alone, it still falls short of GMNScore alone. We believe this is because RelPM, although based on ROT to capture richer structural semantics, still relies on partial matching algorithm, which struggles to fully capture deep semantic equivalence. By contrast, GMNScore directly leverages graph matching to assess functional equivalence, demonstrating strong robustness and stability.
It is worth noting, however, that the effectiveness of combining multiple outcome rewards is significantly lower than that of our proposed Stepwise Reward model, StepRTM. The highest value achieved by the outcome rewards combination is 43.53%, whereas StepRTM achieves a higher value of 45.32%, representing an improvement of 1.79%. This highlights the effectiveness of our designed stepwise reward model StepRTM. More importantly, StepRTM introduces no additional costs in terms of database execution time or GPU memory consumption.
Appendix J Generalization of GMNScore Reward
@@
Although the GMN of GMNScore reward model is trained solely on SQL equivalence data derived from the Spider dataset, it exhibits strong cross-database generalization without requiring any additional fine-tuning. As shown in Table 12, when RL is conducted on the training split of BIRD, GMNScore continues to outperform execution-guided (EX) reward in average TS performance, consistent with the results on training split of Spider.
More broadly, we observe that across both training configurations of Spider and BIRD, our GMNScore consistently achieves higher average validation performance than EX. However, a finer-grained comparison reveals a nuanced trend: on both DeepSeek-Coder-1.3B and 6.7B with PPO, GMNScore consistently outperforms EX on the Spider evaluation dataset but shows a slight performance drop on the BIRD evaluation dataset.
We attribute this discrepancy to the inherent limitations of the GMN model itself. Although we enhanced GMN through data augmentation based on the Spider-train-pair dataset Zhan et al. (2025), the augmented data is still constructed from Spider-style SQL, which biases the model toward better alignment with the structural and stylistic patterns of Spider. In contrast, the BIRD dataset is significantly more challenging, containing more complex SQL constructs such as deeply nested subqueries. As a result, GMN may struggle to judge equivalence in BIRD-style queries as accurately as its performance in Spider. For example, GMN is likely to score generated SQL in Spider-like styles more accurately, leading to more consistent reinforcement signals, while its scoring fidelity for BIRD-like styles is relatively weaker.
In contrast, the EX reward is purely execution-based and thus remains more robust to variations in dataset complexity and style. However, EX also suffers from false positives. Despite this slight drop in the BIRD dataset, GMNScore offers substantially larger improvements on Spider and achieves the best overall average performance. We further hypothesize that stronger base models can compensate for the small drop in reward signal quality in the BIRD dataset.
As discussed previously, GMNScore exhibits slightly reduced BIRD performance compared to EX when using Deepseek-Coder-1.3b/6.7b-Ins models, likely due to GMN’s limited generalization on BIBD-style SQL. However, this performance gap decreases with stronger base models. As shown in Table 13, when using Qwen-Coder-7B and 14B as the policy backbone, GMNScore not only consistently outperforms EX on Spider, but also achieves superior results on BIRD. This further supports the effectiveness and scalability of our reward model framework across different model capacities.
@@
Appendix K Detailed Effectiveness of GMNScore Across GRPO
@@
[…]
@@
[…]
Table 13: TS Performance of Qwen2.5-Coder-7B/14B-Ins models directly trained by PPO/GRPO under EX and GMNScore outcome rewards.
As shown in Table 13, our proposed GMNScore consistently outperforms the EX across both PPO and GRPO training paradigms, and for both Qwen2.5-Coder-7B-Ins and 14B-Ins models. On the 7B model, GMNScore improves the average score from 52.13% to 53.25% under PPO training (+1.12%), and from 52.45% to 54.23% under GRPO training (+1.78%). Similarly, on the 14B model, GMNScore boosts the average score by +1.72% under PPO and +0.62% under GRPO. These consistent improvements demonstrate that GMNScore provides a stable reward signal than EX. Overall, these results confirm the superiority of GMNScore in guiding RL-based Text-to-SQL tasks more effectively than execution-based EX reward model.
Appendix L StepRTM Calculation
@@
Appendix M Training Details of Neural Reward Model with Language odeling
@@
To train the reward model in the above preference pairs, we finetune the pretrained model DeepSeek-Coder-1.3B-Instruct with a reward head. Training is carried out using an effective batch size of $16$ with gradient accumulation $2$. We use the AdamW optimizer, a learning rate of $1.41\times 10^{-5}$, and a linear warm-up schedule. The model is trained for five epochs with a sequence length of 2048 tokens.
@@
The reward model should effectively evaluates the generated SQL based on both syntactic clarity and execution-level correctness. To reinforce schema understanding and functional equivalence, the data was structured as follows:
@@
Appendix N RelNode-based Partial Matching
@@
Based on RelNode representations, a partial matching strategy is employed to measure the similarity between the relational operator trees of the reference SQL $R(q^{\star})$ and the generated SQL R(q^)R($\hat${q}). Let $\mathcal{N}_{q^{\star}}$ and $\mathcal{N}_{\hat{q}}$ denote the sets of nodes in the RelNodes of the reference and generated queries, respectively. A node pair $(n,n^{\prime})$ is considered a match, denoted as $\text{match}(n,n^{\prime})$, if the two nodes share the same relational operator type and their associated attributes are semantically equivalent.
@@
Precision $=\frac{|\{n\in\mathcal{N}_{\hat{q}}\mid\exists n^{\prime}\in\mathcal{N}_{q^{\star}},\ \text{match}(n,n^{\prime})\}|}{|\mathcal{N}_{\hat{q}}|}$
Recall $=\frac{|\{n\in\mathcal{N}_{q^{\star}}\mid\exists n^{\prime}\in\mathcal{N}_{\hat{q}},\ \text{match}(n,n^{\prime})\}|}{|\mathcal{N}_{q^{\star}}|}$
@@
$\begin{aligned} r_{v}^{(t)}&=\mathrm{MLP}(h^{(t)}_{v}\oplus p_{v}^{(t)}),\\ \mu_{v}^{(t+1)}&=\sum_{u\in\mathcal{G}_{2}(v)}a_{u\to v}(r_{v}^{(t)}-r_{u}^{(t)}),\end{aligned}$
@@
Appendix P Prompt Strategy of CTE